OpenStudy (anonymous):

The length of a rectangular park is 6m longer than the width. If the area of the park is 135m^2, find the length and width. it's all about problem solving. anyone help?

OpenStudy (radar):

The area of a rectangle A is equal to length (l) times width (w) or: A=lw let x = width (w) then l (l) is (x + 6)m A= x (x+6) = 135 m^2 x^2 + 6x = 135 m^2 x^2 + 6x - 135 = 0 Solve this quadratic for x.

OpenStudy (radar):

Factor (x + 15)(x - 9)=0 x=9 and x+6= 15 l=15 w=9

OpenStudy (anonymous):

thank you so much.

OpenStudy (radar):

You're welcome, good luck with your studies.

OpenStudy (anonymous):

Thank you so much! :) God bless you & take care! :)

OpenStudy (radar):

Same to you.